1. Find : (i) 641/2 (ii) 321/5 3 (iii) 1251/3

1. Find :
(i) 641/2 (ii) 321/5 3 (iii) 1251/3

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) 641/2 = (82)1/2 = (8)2 #1/2 = 8 (ii) 321/5 = (25)1/5 = (2)5 #1/5 = 2 (iii) 1251/3 = (53)1/3 = (5)3 #1/3 = 5